# Textual semantics and machine learning methods for data product pricing

### v1

@@ 4.3 Classification results based on different feature representation methods
In addition to predicting the exact prices of data products, we also conduct price interval prediction. Specifically, the prices are divided into four quartiles, resulting in five distinct ranges from high to low. After categorizing the prices, the corresponding price ranges are defined as follows: Class 0: 208.33; Class 1: 212.5–416.67; Class 2: 418.5–1,250.0; Class 3: 1,258.33–3,175.0; Class 4: 3,200.0. Such interval-based prediction can provide data providers and data consumers with an approximate price range for reference. In the classification task, we compare the predictive capabilities of different text representation methods (BOW, TF-IDF, Word2Vec, LDA, and BERTopic) in forecasting data product prices. Several prediction models are employed for this experiment, including LR, ANN, DT, SVM, RF, and XGBoost. Table 5 presents the classification prediction results obtained using different text representation methods. In the Table, MR represents the mean results across the six predictive models, Rank indicates the ranking results based on the MR values, and the values in bold font indicate the best results among different text representation methods under the same prediction model.
| Method | LR | ANN | DT | SVM | RF | XGBoost | MR | Rank |
@@ 4.4.1 Regression results
As demonstrated by the results in the figure, the word embeddings generated via text vectorization emerged as the most influential features. Among the top twenty features in the importance ranking, all were word embedding features except for ”industry: Public Sector Data.”
@@
For instance, words with high values along the ’embedding_30’ dimension include: ”demographics, gender, medication, therapeutic, age, …”. These expressions are clearly related to population and health, suggesting that data related to healthcare and individual profiles are generally priced higher. Conversely, words with low values in this dimension include: ”ice, weather, days, …”, indicating that data associated with weather or the environment are generally priced lower.
@@ 4.4.2 Classification results
The differences are notable: the feature embedding_1, which is not significant for the regression model, ranks highest in importance for the classification model. Through back-tracing, we determined that embedding_1 primarily represents macro-level data product descriptions, using terms such as ”national” or ”markets”. Consequently, this feature makes a substantial contribution to classifying the lower price category (Class 1).
@@ 5.2 Practical implications
Based on the XGBoost–mRMR model, we find that model performance does not continue to improve as the number of features increases. Instead, optimal or near-optimal prediction performance is achieved with roughly 30 features for the classification task and around 60 features for the regression task. This result suggests that the value of data products is largely concentrated in a limited set of high-information features. For data providers, simply adding more attributes when submitting products does not necessarily raise the price; rather, retaining a concise set of high-information-value features is more beneficial. For data consumers, it is unnecessary to examine every listed attribute in detail; attention should instead focus on the key determinants of value. For data platform, feature-selection techniques such as mRMR can be incorporated into pricing assistance tools to enhance platform efficiency and improve matching quality in data trade.

### v2

@@ 4.3 Classification results based on different feature representation methods
In the classification task, we compare the predictive capabilities of different text representation methods (BOW, TF-IDF, Word2Vec, LDA, and BERTopic) in forecasting data product prices. Several prediction models are employed for this experiment, including LR, ANN, DT, SVM, RF, and XGBoost. Table 6 presents the classification prediction results obtained using different text representation methods. In the Table, MR represents the mean results across the six predictive models, Rank indicates the ranking results based on the MR values, and the values in bold font indicate the best results among different text representation methods under the same prediction model.
| Method | LR | ANN | DT | SVM | RF | XGBoost | MV | Rank |
@@
The performance differences between Word2Vec and BOW/TF-IDF can be explained by the fundamental distinctions between regression-based and classification-based pricing tasks. Regression tasks aim to model continuous variations in data product prices and therefore require the model to capture gradual changes in semantic intensity, enabling a smooth mapping between textual semantics and price values. In the context of data product pricing, prices are rarely determined by a single keyword; instead, they emerge from the joint effects of multiple latent semantic dimensions, such as functional richness, data coverage, application scenarios, update frequency, and timeliness. By contrast, classification tasks focus on predicting discrete price levels. Their primary objective is to separate samples into a limited number of categories, emphasizing the identification of clear decision boundaries rather than modeling continuous semantic variation. In such settings, the presence of a small number of highly discriminative keywords is often sufficient to distinguish price levels effectively.
From a representation perspective, Word2Vec maps textual content into a continuous, low-dimensional, and dense semantic space, which makes it particularly well suited for regression tasks. First, data product prices are continuous variables, and the continuity of the Word2Vec semantic space aligns well with this property, facilitating the learning of smooth functional relationships in regression models. Second, semantically similar words are embedded in nearby regions of the vector space, allowing the model to capture the cumulative effect of semantic intensity on pricing while reducing sensitivity to lexical differences. In contrast, BOW and TF-IDF generate high-dimensional and sparse representations that ignore contextual information and semantic similarity, focusing instead on word occurrence and frequency. These characteristics are advantageous in classification tasks, as discriminative keywords linked to specific price tiers are preserved with minimal noise.
@@ 4.4.1 Regression results
As demonstrated by the results in the figure, the word embeddings generated via text vectorization emerged as the most influential features. Among the top twenty features in the importance ranking, all were word embedding features except for “industry: Public Sector Data.”
@@
For instance, words with high values along the ‘embedding_30’ dimension include: “demographics, gender, medication, therapeutic, age, …”. These expressions are clearly related to population and health, suggesting that data related to healthcare and individual profiles are generally priced higher. Conversely, words with low values in this dimension include: “ice, weather, days, …”, indicating that data associated with weather or the environment are generally priced lower.
@@ 4.4.2 Classification results
The differences are notable: the feature embedding_1, which is not significant for the regression model, ranks highest in importance for the classification model. Through back-tracing, we determined that embedding_1 primarily represents macro-level data product descriptions, using terms such as “national” or “markets”. Consequently, this feature makes a substantial contribution to classifying the lower price category (Class 1).
@@ 5.2 Practical implications
Based on the XGBoost-mRMR model, we find that model performance does not continue to improve as the number of features increases. Instead, optimal or near-optimal prediction performance is achieved with roughly 30 features for the classification task and around 60 features for the regression task. This result suggests that the value of data products is largely concentrated in a limited set of high-information features. For data providers, simply adding more attributes when submitting products does not necessarily raise the price; rather, retaining a concise set of high-information-value features is more beneficial. For data consumers, it is unnecessary to examine every listed attribute in detail; attention should instead focus on the key determinants of value. For data platform, feature-selection techniques such as mRMR can be incorporated into pricing assistance tools to enhance platform efficiency and improve matching quality in data trade.
@@ 6 Conclusion
Acknowledgments
The authors acknowledge the financial support from the National Natural Science Foundation of China (72192802, 72342008, 72401029 and 72473013).
Declaration of Competing Interests
The authors declare no competing interests.
Appendix
Appendix A. Prompts